# Vision-Based Mistake Analysis in Procedural Activities: A Review of Advances and Challenges

### v1

@@ I Introduction
Making mistakes is intrinsic to human nature. While they can act as critical drivers for learning and skill refinement [1], they may also lead to costly or dangerous consequences in high-stakes environments. The ability to detect, anticipate, and respond to mistakes, ideally before they escalate into failures, is essential in contexts where safety, efficiency, or performance are paramount. In such scenarios, real-time mistake analysis can support timely interventions and corrective actions, ultimately improving both human and system-level outcomes.
Figure 1: Illustration of mistake analysis in an industrial assembly task: (1) recognize an unexpected or omitted action given an activity protocol (mistake recognition), e.g.,the worker “attached part C” instead of “grabbing part B”, (2) temporally detect an error within the execution of an action (mistake detection) and (3) predict a mistake before it fully occurs (early mistake recognition), e.g., warn the worker about selecting the wrong tool. The conceptual worker sketch was generated with [2].
@@
The impact of robust mistake analysis methods spans multiple domains. In industrial automation, they can enhance productivity by minimizing human error, ensuring proper task execution, and reducing the risk of injuries (see Figure 1). In human-robot collaboration, anticipating potential mistakes facilitates smoother cooperation, reduces failure rates, and improves task handover efficiency. Similarly, in rehabilitation, education, or assisted living, such systems can guide users through complex tasks, offering feedback and corrections that enhance skill acquisition or daily independence. In embodied robotics, mistake analysis plays a dual role: supporting human–robot interaction [3] and enabling robots to recognize and recover from abnormal or unexpected situations [4], a critical step toward robust autonomy in real-world environments
@@
Computer vision has emerged as a powerful modality for mistake detection in procedural tasks. Unlike other sensor-based methods, vision offers a non-invasive, cost-effective, and scalable alternative leveraging rich visual data. Advances in activity recognition, action detection, and multimodal understanding [5] enable detecting subtle deviations such as incorrect motion trajectories, missing steps, or tool misuse. For instance, in cooking, a vision system might infer an error when a user prematurely switches tools or leaves a pan unattended.
Despite recent progress, mistake analysis in procedural activities remains highly challenging [6, 7]. A major difficulty lies in the variability of task execution: even well-defined procedures often allow multiple valid action sequences rather than a single canonical order. This complicates the distinction between genuine errors and acceptable alternatives. Further ambiguity arises at the action level, where steps vary in motion dynamics, timing, or appearance, making the boundary between atypical yet correct and truly erroneous behavior non-trivial. Additional challenges stem from viewpoint changes (e.g., egocentric vs. exocentric), object multifunctionality, and activity compositionality, which demand fine-grained spatiotemporal reasoning. Crucially, mistake detection requires judging whether an action is appropriate in context, making it a higher-order reasoning task beyond static recognition.
@@ II Mistake Analysis: Problem Statement
Figure 3: Illustration of the association between video anomaly and mistake detection in procedural activities.Human sketches in the figure generated with [8].
@@ II-A Distinguishing Mistake Analysis in Procedural Activities from Conventional Video Anomaly Detection
In principle, mistake detection in procedural activities is related to video anomaly detection [9, 10]. However, the two tasks differ fundamentally in the contextual grounding of what constitutes an error, with mistake detection requiring task-specific procedural knowledge to distinguish between acceptable variations and true deviations (Figure 3).
Video anomaly detection typically focuses on identifying irregularities, i.e., deviations from normal visual or motion patterns, without necessarily understanding task-specific goals. Such anomalies exhibit distinct visual or semantic properties, making them recognizable as unexpected events. For instance, in the ShanghaiTech Campus dataset [11], a widely used benchmark, anomalies include unusual behaviors such as fighting, chasing, or other atypical activities, detected through their visual distinctiveness from typical campus scenes. Procedural mistake detection differs markedly from conventional anomaly detection. Unlike generic anomalies, mistakes in procedural activities are defined relative to a structured sequence of actions required to complete a task. These include missing, adding, modifying, or incorrectly executing an action step, all of which must be interpreted within the overarching goal. For example, an engineer assembling a circuit board who omits a resistor must disassemble parts to correct the mistake. Such errors stem from violating the prescribed action order and cannot be captured by appearance-based irregularities alone but require reasoning over long temporal dependencies. In contrast, video anomaly detection primarily flags unusual events based on appearance or motion, whereas mistake detection is inherently structured and goal-dependent.
@@
As a final remark, most anomaly detection methods operate on static or short-term temporal observations, whereas mistake detection in procedural activities demands long-range temporal reasoning to assess whether an action sequence aligns with a predefined procedure. This challenge is further amplified in egocentric video [12, 13, 14], where shifting viewpoints, occlusions and object-scale variations hinder the application of conventional anomaly detection techniques.
@@ III A Taxonomy of Mistakes
Due to the emerging nature of the problem, however, existing works adopt varied definitions of mistake/error types, particularly for executional mistakes. For instance, Peddi et al. [13], in the CaptainCook4D dataset, define the categories {Order Error, Omission Error, Technique Error, Timing Error, Temperature Error}, with the first two being procedural and the remainder executional. In contrast, Lee et al. [15] propose a different but partially overlapping scheme, distinguishing Step Omission, Step Addition and Step Correction as procedural, while identifying Step Modification and Step Slip as executional. Although both taxonomies include omissions and technique-based deviations, their definitions differ in granularity and interpretation. A general overview of the taxonomy adopted in this survey is illustrated in Figure 4. Following the distinctions introduced in prior work [16], procedural mistakes can be further divided into the following cases:
@@ IV Mistake Analysis Datasets
Figure 5: Timeline of datasets supporting mistake analysis (2018–2025). Circle sizes refer to dataset scale. Datasets: Epic-Tent [17], CSV [18], Assembly-101 [14], BRIO-TA [19], ATA [20], HoloAssist [21], CaptainCook4D [13], Ego-Exo4D [22], IndustReal [23], EgoPER [15] and EgoOops [12].
@@ IV-A Datasets: overview & key attributes
Epic-Tent [17]: Was the first dataset to offer annotated task-level mistakes in a real-world procedural activity setting. It comprises over 5.4 hours of first-person (ego-centric) video recorded from 29 participants, each equipped with two head-mounted cameras, tasked with assembling a tent in an outdoor setting. Epic-Tent includes rich multi-modal annotations, such as frame-level action and error labels, 2D gaze positions, and self-reported uncertainty levels, enabling a range of research applications in egocentric action and activity recognition, human-object interaction, and mistake detection. It supports coarse-to-fine procedural understanding, with the overall tent assembly task decomposed into 12 sub-activities, each comprising of low-level steps from a total of 38 unique actions.
@@
For mistake analysis, Epic-Tent provides coarse, segment-level annotations indicating whether the overall task or its sub-components were completed correctly or incorrectly, without distinguishing between procedural (e.g., skipping a step, incorrect order) and executional (e.g., misplacing a pole, using incorrect force) errors, though both may occur. Unlike other datasets targeting mistake analysis, Epic-Tent did not define evaluation protocols for this task, as nearly every sample included at least one procedural error, preventing binary success/failure splits. To address these limitations and explicitly enable online mistake detection, Flaborea et al. [24] introduced Epic-Tent-O, proposing an uncertainty-based dataset partitioning strategy: videos from high-confidence participants were used for training, while those from less confident performers, more prone to errors, were reserved for testing.
@@
Chemical Sequence Verification (CSV) [18]: CSV addresses the challenges of procedural activity understanding and mistake analysis in the context of chemical procedures. It encompasses 14 activities, each representing a chemical experiment conducted by 82 volunteers following predefined scripts. The dataset comprises a collection of 70 video recordings, partitioned into train-test subsets under a 80-20 split. The recorded sequences include deviations from the predefined execution protocols, incorporating step-level transformations such as additions, deletions and reordering of procedural steps. Annotations follow a weak supervision scheme, providing video-level activity and action labels without temporal boundaries for the actions. Notably, the dataset does not include explicit annotations specifying the procedural mistake types. The dataset supports multiple tasks, including activity, action and mistake recognition. Additionally, it introduces the task of early mistake recognition (early warning) to predict errors before they fully manifest, enabling proactive intervention.
@@
Assembly101 [14]: This is a large-scale benchmark designed for action and activity understanding in assembly tasks, with a focus on fine-grained procedural activities. It contains over 362 videos which amount for 6,000 video clips of individuals assembling 101 toy car models, captured from multiple cameras (exocentric and egocentric) and viewing angles in a constrained laboratory environment. Assembly101 supports a wide range of tasks, including fine-grained action recognition, action localization, action anticipation, activity recognition and mistake or error detection. Additionally, the dataset is designed to enable evaluation in zero-shot settings, allowing models to generalize to previously unseen actions or activity compositions. Each video is annotated with frame-level action annotations under two granularity levels (1380 fine- (single-task) and 202 coarse-grained (multiple fine-grained)) and mistake annotations, making it suitable for both step-level and task-level reasoning. The dataset provides multi-modal data, including synchronized RGB video from multiple viewpoints, depth information and 3D hand pose data.
@@
For mistake analysis tasks, annotations exist for 1.01K of the 6K benchmark video clips, provided at a coarse action-segment level, focusing on procedural errors, such as missing, or out-of-order steps. While executional errors (e.g., incorrect orientation or grasp) may occur in the data, they are relatively rare and not annotated. The annotation process does not differentiate between procedural and executional errors and no formal taxonomy of the mistake types is provided to support fine-grained mistake recognition. Subsequent works by Ding et al. [16] and Flabora et al. [24] extended Assembly101 to better support the mistake detection task with additional annotations and better restructuring. Specifically, Ding et al. [16] enriched Assembly101 with mistake-type information and explicit part-to-part connection details to facilitate procedural and mistake reasoning, whereas Flabora et al. [24] enabled the support of online mistake detection, producing the variant Assembly101-O. In this adaptation, all correctly executed sequences were moved to the training set, while sequences with mistakes were reserved for validation and testing. Procedure lengths were adjusted to better suit real-time detection requirements.
@@
BRIO Toy Assembly (BRIO-TA) [19]: The dataset comprises 75 video recordings capturing both normal and anomalous assembly processes of a toy car model. The dataset includes three distinct error types that may arise during the assembly process: (a) step ordering variation, (b) step omission and (c) abnormal duration. Each erroneous sequence exhibits only one of these three anomaly types. The dataset provides segment-wise temporal annotations of action occurrences along with video-level mistake annotations. It supports the tasks of action recognition and segmentation, as well as mistake recognition. RGB videos is the only supported modality. Performance is evaluated using standard action segmentation metrics, including Accuracy, Intersection over Union (IoU) and Edit Distance, alongside mistake recognition metrics such as Accuracy. However, the dataset does not explicitly define an evaluation protocol or metrics tailored to each specific error type.
@@
Anomalous Toy Assembly (ATA) [20]: The design of the ATA dataset was motivated by the challenge of identifying mistakes in procedural activities, where the absence of fine-grained temporal annotations poses significant difficulties. The dataset comprises of 141 untrimmed videos of 32 participants assembling 3 different toy models, recorded from 4 distinct viewpoints. It includes both correct and erroneous executions, where errors can be seen or unseen during training, making it suitable for evaluating generalization in mistake detection. ATA is annotated with weak supervision in the form of video transcripts describing action sequences, while errors are not explicitly labeled in the transcripts, simulating real-world conditions where mistake annotations are unavailable. The captured errors encompass execution mistakes (e.g., misplacement of parts) and procedural mistakes (e.g., step omissions or order deviations), making it a comprehensive benchmark for evaluating mistake detection methods for both supervised and zero-shot settings. The dataset provides multiple modalities, including RGB videos, ASR (automatic speech recognition) transcripts and gaze data, enabling multimodal learning.
@@
HoloAssist [21]: The HoloAssist dataset was designed to advance interactive AI assistants in real-world settings. It captures 20 collaborative object-centric manipulation tasks between two individuals: a task performer, who executes the task while wearing a mixed-reality headset that records seven synchronized data streams, and a task instructor, who observes the performer’s first-person feed in real time and provides verbal guidance, including interventions in case of mistakes. The dataset includes 166 hours of interaction data from 350 instructor–performer pairs drawn from 222 participants. Manual annotations accompany the dataset, including text summaries, intervention types, mistake labels, and segment-level action annotations at two abstraction levels: (a) coarse-grained, describing high-level task steps, and (b) fine-grained, referring to atomic actions composing each step. There are 414 coarse-grained and 1,887 fine-grained action classes.
@@
CaptainCook4D [13]: This is one of the largest and most well structured datasets targeting the understanding of mistakes/errors in complex procedural activities within the domain of cooking. It comprises 384 recordings totaling over 94.5 hours of real-world kitchen interactions, capturing both standard and intentionally erroneous executions of 24 recipes. The activities are structured from a pool of 352 coarse-grained actions which are, in turn, defined from a set of fine-grained actions (steps). A standout feature of CaptainCook4D is its multimodal design, which includes RGB video, depth data, 3D object annotations, gaze data and audio. Unlike previous datasets that either lacked error annotations or only considered a binary notion of correctness, CaptainCook4D offers a rich and fine-grained taxonomy of 7 mistake types that capture both procedural and executional deviations. These mistakes are annotated at both the coarse- and fine-grained levels, allowing models to reason not only about what went wrong but also when and how this occurred.
@@
Ego-Exo4D [25]: This dataset was designed to support research in ego-exo video learning and multi-modal perception, focusing on skilled human activities. It is among the largest publicly available datasets containing time-synchronized first- and third-person video data, collected from 740 participants across 123 scenes in 13 cities worldwide. Participants performed 43 skilled physical and procedural activities unscripted within natural settings. In addition to RGB video, the dataset provides multi-modal sensory data, including audio, IMU, eye gaze, RGB and grayscale SLAM recordings, and 3D environment point clouds. It also includes time-indexed video-language resources, such as participant narrations during task execution and spoken expert commentaries evaluating performance. The dataset supports tasks related to activity understanding, including modeling instructor-learner relationships, recognizing fine-grained key-steps and task structures, assessing skill proficiency, and recovering 3D body and hand movements from egocentric video. Apart from action(step) and activity annotations, it provides 2D/3D pose annotations and object annotations, with an average of 5.5 objects annotated with correspondences between the two views in each take. Mistake analysis is introduced as a downstream problem within the task of procedure understanding. For this purpose, only 6 activities are considered, covering 186 step annotations. In total 628 sequences are recorded. The focus is on detecting procedural mistakes by identifying missing steps and recognizing steps that were not intended for the given activity.
@@
IndustReal [23]: This dataset was created to support the task of procedure step recognition in the context of industrial settings, with a particular focus on handling execution errors. The dataset contains 84 egocentric (first-person) videos, from 27 participants, captured during the assembly of a toy car, where the primary goal is to recognize and understand individual steps of the assembly process and detect execution errors within these steps. IndustReal focuses on two activities (a) the activity of assembling a toy car and (b) the activity of maintenance. The first is the most challenging, with the task being divided into 5 sub-activities, each containing multiple actions (steps). The total set of distinct actions in the dataset is 75. The structured breakdown of the assembly activity enables the tracking of task progression by identifying which sub-activity (or sub-task) has been completed. The dataset is enriched with multimodal sensor data, including video recordings and depth information (from stereo or depth cameras), providing a richer understanding of spatial and temporal context.
@@
EgoPER [15]: EgoPER is a recent dataset that introduces a multi-modal benchmark for the study of mistake/error detection in egocentric procedural tasks. The dataset features 5 diverse cooking activities preparation of pinwheels, coffee, quesadilla, tea and oatmeal performed by 11 participants. It contains over 120 hours of video recordings captured with a head-mounted camera in unconstrained environments. Unlike prior datasets that focus solely on successful task completions and generalized mistake recognition, EgoPER explicitly incorporates a rich taxonomy of human errors, including step omission, addition, modification, slip, and correction. Although all mistakes/errors are scripted, the execution of those errors was left natural, meaning participants had flexibility in how they introduced the error, leading to realistic and diverse instantiations of each error type. Each video is annotated with segment-wise action labels, object bounding boxes, hand-object interactions, gaze tracking, and audio, offering a comprehensive view of both correct and erroneous executions.

### v2

@@ I Introduction
Making mistakes is intrinsic to human nature. While they can act as critical drivers for learning and skill refinement [reason1990human], they may also lead to costly or dangerous consequences in high-stakes environments. The ability to detect, anticipate, and respond to mistakes, ideally before they escalate into failures, is essential in contexts where safety, efficiency, or performance are paramount. In such scenarios, real-time mistake analysis can support timely interventions and corrective actions, ultimately improving both human and system-level outcomes.
Figure 1: Illustration of mistake analysis in an industrial assembly task: (1) recognize an unexpected or omitted action given an activity protocol (mistake recognition), e.g.,the worker “attached part C” instead of “grabbing part B”, (2) temporally detect an error within the execution of an action (mistake detection) and (3) predict a mistake before it fully occurs (early mistake recognition), e.g., warn the worker about selecting the wrong tool. The conceptual worker sketch was generated with [google_gemini].
@@
The impact of robust mistake analysis methods spans multiple domains. In industrial automation, they can enhance productivity by minimizing human error, ensuring proper task execution, and reducing the risk of injuries (see Figure 1). In human-robot collaboration, anticipating potential mistakes facilitates smoother cooperation, reduces failure rates, and improves task handover efficiency. Similarly, in rehabilitation, education, or assisted living, such systems can guide users through complex tasks, offering feedback and corrections that enhance skill acquisition or daily independence. In embodied robotics, mistake analysis plays a dual role: supporting human–robot interaction [nikolaidis2017human] and enabling robots to recognize and recover from abnormal or unexpected situations [garrabe2024enhancing], a critical step toward robust autonomy in real-world environments
@@
Computer vision has emerged as a powerful modality for mistake detection in procedural tasks. Unlike other sensor-based methods, vision offers a non-invasive, cost-effective, and scalable alternative leveraging rich visual data. Advances in activity recognition, action detection, and multimodal understanding [stergiou2025time] enable detecting subtle deviations such as incorrect motion trajectories, missing steps, or tool misuse. For instance, in cooking, a vision system might infer an error when a user prematurely switches tools or leaves a pan unattended.
Despite recent progress, mistake analysis in procedural activities remains highly challenging [aggarwal2011human, herath2017going]. A major difficulty lies in the variability of task execution: even well-defined procedures often allow multiple valid action sequences rather than a single canonical order. This complicates the distinction between genuine errors and acceptable alternatives. Further ambiguity arises at the action level, where steps vary in motion dynamics, timing, or appearance, making the boundary between atypical yet correct and truly erroneous behavior non-trivial. Additional challenges stem from viewpoint changes (e.g., egocentric vs. exocentric), object multifunctionality, and activity compositionality, which demand fine-grained spatiotemporal reasoning. Crucially, mistake detection requires judging whether an action is appropriate in context, making it a higher-order reasoning task beyond static recognition.
@@ II Mistake Analysis: Problem Statement
Figure 3: Illustration of the association between video anomaly and mistake detection in procedural activities.Human sketches in the figure generated with [openai_chatgpt].
@@ II-A Distinguishing Mistake Analysis in Procedural Activities from Conventional Video Anomaly Detection
In principle, mistake detection in procedural activities is related to video anomaly detection [V_Anomaly1, V_Anomaly2]. However, the two tasks differ fundamentally in the contextual grounding of what constitutes an error, with mistake detection requiring task-specific procedural knowledge to distinguish between acceptable variations and true deviations (Figure 3).
Video anomaly detection typically focuses on identifying irregularities, i.e., deviations from normal visual or motion patterns, without necessarily understanding task-specific goals. Such anomalies exhibit distinct visual or semantic properties, making them recognizable as unexpected events. For instance, in the ShanghaiTech Campus dataset [ShanghaiTech], a widely used benchmark, anomalies include unusual behaviors such as fighting, chasing, or other atypical activities, detected through their visual distinctiveness from typical campus scenes. Procedural mistake detection differs markedly from conventional anomaly detection. Unlike generic anomalies, mistakes in procedural activities are defined relative to a structured sequence of actions required to complete a task. These include missing, adding, modifying, or incorrectly executing an action step, all of which must be interpreted within the overarching goal. For example, an engineer assembling a circuit board who omits a resistor must disassemble parts to correct the mistake. Such errors stem from violating the prescribed action order and cannot be captured by appearance-based irregularities alone but require reasoning over long temporal dependencies. In contrast, video anomaly detection primarily flags unusual events based on appearance or motion, whereas mistake detection is inherently structured and goal-dependent.
@@
As a final remark, most anomaly detection methods operate on static or short-term temporal observations, whereas mistake detection in procedural activities demands long-range temporal reasoning to assess whether an action sequence aligns with a predefined procedure. This challenge is further amplified in egocentric video [EgoOops, CaptainCook4D, Assembly101], where shifting viewpoints, occlusions and object-scale variations hinder the application of conventional anomaly detection techniques.
@@ III A Taxonomy of Mistakes
Due to the emerging nature of the problem, however, existing works adopt varied definitions of mistake/error types, particularly for executional mistakes. For instance, Peddi et al. [CaptainCook4D], in the CaptainCook4D dataset, define the categories {Order Error, Omission Error, Technique Error, Timing Error, Temperature Error}, with the first two being procedural and the remainder executional. In contrast, Lee et al. [EgoPED] propose a different but partially overlapping scheme, distinguishing Step Omission, Step Addition and Step Correction as procedural, while identifying Step Modification and Step Slip as executional. Although both taxonomies include omissions and technique-based deviations, their definitions differ in granularity and interpretation. A general overview of the taxonomy adopted in this survey is illustrated in Figure 4. Following the distinctions introduced in prior work [ding2023mistakecountsassembly], procedural mistakes can be further divided into the following cases:
@@ IV Mistake Analysis Datasets
Figure 5: Timeline of datasets supporting mistake analysis (2018–2025). Circle sizes refer to dataset scale. Datasets: Epic-Tent [EPICTENT], CSV [CSV], Assembly-101 [Assembly101], BRIO-TA [BRIOTA], ATA [ATA], HoloAssist [HoloAssist], CaptainCook4D [CaptainCook4D], Ego-Exo4D [grauman2022ego4d], IndustReal [IndustReal], EgoPER [EgoPED] and EgoOops [EgoOops]. Ego4D-M and EK-M [li2025mistake] are synthetic extensions of Ego4D [grauman2022ego4d] and EK-100 [damen2022rescaling], generated to incorporate explicit mistake annotations.
@@ IV-A Datasets: overview & key attributes
Epic-Tent [EPICTENT]: Was the first dataset to offer annotated task-level mistakes in a real-world procedural activity setting. It comprises over 5.4 hours of first-person (ego-centric) video recorded from 29 participants, each equipped with two head-mounted cameras, tasked with assembling a tent in an outdoor setting. Epic-Tent includes rich multi-modal annotations, such as frame-level action and error labels, 2D gaze positions, and self-reported uncertainty levels, enabling a range of research applications in egocentric action and activity recognition, human-object interaction, and mistake detection. It supports coarse-to-fine procedural understanding, with the overall tent assembly task decomposed into 12 sub-activities, each comprising of low-level steps from a total of 38 unique actions.
@@
For mistake analysis, Epic-Tent provides coarse, segment-level annotations indicating whether the overall task or its sub-components were completed correctly or incorrectly, without distinguishing between procedural (e.g., skipping a step, incorrect order) and executional (e.g., misplacing a pole, using incorrect force) errors, though both may occur. Unlike other datasets targeting mistake analysis, Epic-Tent did not define evaluation protocols for this task, as nearly every sample included at least one procedural error, preventing binary success/failure splits. To address these limitations and explicitly enable online mistake detection, Flaborea et al. [flaborea2024prego] introduced Epic-Tent-O, proposing an uncertainty-based dataset partitioning strategy: videos from high-confidence participants were used for training, while those from less confident performers, more prone to errors, were reserved for testing.
@@
Chemical Sequence Verification (CSV) [CSV]: CSV addresses the challenges of procedural activity understanding and mistake analysis in the context of chemical procedures. It encompasses 14 activities, each representing a chemical experiment conducted by 82 volunteers following predefined scripts. The dataset comprises a collection of 70 video recordings, partitioned into train-test subsets under a 80-20 split. The recorded sequences include deviations from the predefined execution protocols, incorporating step-level transformations such as additions, deletions and reordering of procedural steps. Annotations follow a weak supervision scheme, providing video-level activity and action labels without temporal boundaries for the actions. Notably, the dataset does not include explicit annotations specifying the procedural mistake types. The dataset supports multiple tasks, including activity, action and mistake recognition. Additionally, it introduces the task of early mistake recognition (early warning) to predict errors before they fully manifest, enabling proactive intervention.
@@
Assembly101 [Assembly101]: This is a large-scale benchmark designed for action and activity understanding in assembly tasks, with a focus on fine-grained procedural activities. It contains over 362 videos which amount for 6,000 video clips of individuals assembling 101 toy car models, captured from multiple cameras (exocentric and egocentric) and viewing angles in a constrained laboratory environment. Assembly101 supports a wide range of tasks, including fine-grained action recognition, action localization, action anticipation, activity recognition and mistake or error detection. Additionally, the dataset is designed to enable evaluation in zero-shot settings, allowing models to generalize to previously unseen actions or activity compositions. Each video is annotated with frame-level action annotations under two granularity levels (1380 fine- (single-task) and 202 coarse-grained (multiple fine-grained)) and mistake annotations, making it suitable for both step-level and task-level reasoning. The dataset provides multi-modal data, including synchronized RGB video from multiple viewpoints, depth information and 3D hand pose data.
@@
For mistake analysis tasks, annotations exist for 1.01K of the 6K benchmark video clips, provided at a coarse action-segment level, focusing on procedural errors, such as missing, or out-of-order steps. While executional errors (e.g., incorrect orientation or grasp) may occur in the data, they are relatively rare and not annotated. The annotation process does not differentiate between procedural and executional errors and no formal taxonomy of the mistake types is provided to support fine-grained mistake recognition. Subsequent works by Ding et al. [ding2023mistakecountsassembly] and Flabora et al. [flaborea2024prego] extended Assembly101 to better support the mistake detection task with additional annotations and better restructuring. Specifically, Ding et al. [ding2023mistakecountsassembly] enriched Assembly101 with mistake-type information and explicit part-to-part connection details to facilitate procedural and mistake reasoning, whereas Flabora et al. [flaborea2024prego] enabled the support of online mistake detection, producing the variant Assembly101-O. In this adaptation, all correctly executed sequences were moved to the training set, while sequences with mistakes were reserved for validation and testing. Procedure lengths were adjusted to better suit real-time detection requirements.
@@
BRIO Toy Assembly (BRIO-TA) [BRIOTA]: The dataset comprises 75 video recordings capturing both normal and anomalous assembly processes of a toy car model. The dataset includes three distinct error types that may arise during the assembly process: (a) step ordering variation, (b) step omission and (c) abnormal duration. Each erroneous sequence exhibits only one of these three anomaly types. The dataset provides segment-wise temporal annotations of action occurrences along with video-level mistake annotations. It supports the tasks of action recognition and segmentation, as well as mistake recognition. RGB videos is the only supported modality. Performance is evaluated using standard action segmentation metrics, including Accuracy, Intersection over Union (IoU) and Edit Distance, alongside mistake recognition metrics such as Accuracy. However, the dataset does not explicitly define an evaluation protocol or metrics tailored to each specific error type.
@@
Anomalous Toy Assembly (ATA) [ATA]: The design of the ATA dataset was motivated by the challenge of identifying mistakes in procedural activities, where the absence of fine-grained temporal annotations poses significant difficulties. The dataset comprises of 141 untrimmed videos of 32 participants assembling 3 different toy models, recorded from 4 distinct viewpoints. It includes both correct and erroneous executions, where errors can be seen or unseen during training, making it suitable for evaluating generalization in mistake detection. ATA is annotated with weak supervision in the form of video transcripts describing action sequences, while errors are not explicitly labeled in the transcripts, simulating real-world conditions where mistake annotations are unavailable. The captured errors encompass execution mistakes (e.g., misplacement of parts) and procedural mistakes (e.g., step omissions or order deviations), making it a comprehensive benchmark for evaluating mistake detection methods for both supervised and zero-shot settings. The dataset provides multiple modalities, including RGB videos, ASR (automatic speech recognition) transcripts and gaze data, enabling multimodal learning.
@@
HoloAssist [HoloAssist]: The HoloAssist dataset was designed to advance interactive AI assistants in real-world settings. It captures 20 collaborative object-centric manipulation tasks between two individuals: a task performer, who executes the task while wearing a mixed-reality headset that records seven synchronized data streams, and a task instructor, who observes the performer’s first-person feed in real time and provides verbal guidance, including interventions in case of mistakes. The dataset includes 166 hours of interaction data from 350 instructor–performer pairs drawn from 222 participants. Manual annotations accompany the dataset, including text summaries, intervention types, mistake labels, and segment-level action annotations at two abstraction levels: (a) coarse-grained, describing high-level task steps, and (b) fine-grained, referring to atomic actions composing each step. There are 414 coarse-grained and 1,887 fine-grained action classes.
@@
CaptainCook4D [CaptainCook4D]: This is one of the largest and most well structured datasets targeting the understanding of mistakes/errors in complex procedural activities within the domain of cooking. It comprises 384 recordings totaling over 94.5 hours of real-world kitchen interactions, capturing both standard and intentionally erroneous executions of 24 recipes. The activities are structured from a pool of 352 coarse-grained actions which are, in turn, defined from a set of fine-grained actions (steps). A standout feature of CaptainCook4D is its multimodal design, which includes RGB video, depth data, 3D object annotations, gaze data and audio. Unlike previous datasets that either lacked error annotations or only considered a binary notion of correctness, CaptainCook4D offers a rich and fine-grained taxonomy of 7 mistake types that capture both procedural and executional deviations. These mistakes are annotated at both the coarse- and fine-grained levels, allowing models to reason not only about what went wrong but also when and how this occurred.
@@
Ego-Exo4D [grauman2024ego]: This dataset was designed to support research in ego-exo video learning and multi-modal perception, focusing on skilled human activities. It is among the largest publicly available datasets containing time-synchronized first- and third-person video data, collected from 740 participants across 123 scenes in 13 cities worldwide. Participants performed 43 skilled physical and procedural activities unscripted within natural settings. In addition to RGB video, the dataset provides multi-modal sensory data, including audio, IMU, eye gaze, RGB and grayscale SLAM recordings, and 3D environment point clouds. It also includes time-indexed video-language resources, such as participant narrations during task execution and spoken expert commentaries evaluating performance. The dataset supports tasks related to activity understanding, including modeling instructor-learner relationships, recognizing fine-grained key-steps and task structures, assessing skill proficiency, and recovering 3D body and hand movements from egocentric video. Apart from action(step) and activity annotations, it provides 2D/3D pose annotations and object annotations, with an average of 5.5 objects annotated with correspondences between the two views in each take. Mistake analysis is introduced as a downstream problem within the task of procedure understanding. For this purpose, only 6 activities are considered, covering 186 step annotations. In total 628 sequences are recorded. The focus is on detecting procedural mistakes by identifying missing steps and recognizing steps that were not intended for the given activity.
@@
IndustReal [IndustReal]: This dataset was created to support the task of procedure step recognition in the context of industrial settings, with a particular focus on handling execution errors. The dataset contains 84 egocentric (first-person) videos, from 27 participants, captured during the assembly of a toy car, where the primary goal is to recognize and understand individual steps of the assembly process and detect execution errors within these steps. IndustReal focuses on two activities (a) the activity of assembling a toy car and (b) the activity of maintenance. The first is the most challenging, with the task being divided into 5 sub-activities, each containing multiple actions (steps). The total set of distinct actions in the dataset is 75. The structured breakdown of the assembly activity enables the tracking of task progression by identifying which sub-activity (or sub-task) has been completed. The dataset is enriched with multimodal sensor data, including video recordings and depth information (from stereo or depth cameras), providing a richer understanding of spatial and temporal context.
@@
EgoPER [EgoPED]: EgoPER is a recent dataset that introduces a multi-modal benchmark for the study of mistake/error detection in egocentric procedural tasks. The dataset features 5 diverse cooking activities preparation of pinwheels, coffee, quesadilla, tea and oatmeal performed by 11 participants. It contains over 120 hours of video recordings captured with a head-mounted camera in unconstrained environments. Unlike prior datasets that focus solely on successful task completions and generalized mistake recognition, EgoPER explicitly incorporates a rich taxonomy of human errors, including step omission, addition, modification, slip, and correction. Although all mistakes/errors are scripted, the execution of those errors was left natural, meaning participants had flexibility in how they introduced the error, leading to realistic and diverse instantiations of each error type. Each video is annotated with segment-wise action labels, object bounding boxes, hand-object interactions, gaze tracking, and audio, offering a comprehensive view of both correct and erroneous executions.